what is the degree of freedom for a binary azeotrope? explain with examples.

Grade:12th pass

1 Answers

Arun
5 years ago
binarysystem containing an azeotropicmixture in equilibrium with its vapor has two species, two phases, and one relation among intensive variables: \(x\A =y\A\). The number of degrees of freedom is then \(F = 2+s-r-P = 2+2-1-2 = 1\); the system is univariant